Gemini, Inc., an all-equity firm, is considering an investment of $1.74 million that will be depreciated according to the straight-line method over its four-year life. The project is expected to generate earnings before taxes and depreciation of $608,000 per year for four years. The investment will not change the risk level of the firm. The company can obtain a four-year, 8.9 percent loan to finance the project from a local bank. All principal will be repaid in one balloon payment at the end of the fourth year. The bank will charge the firm $58,000 in flotation fees, which will be amortized over the four-year life of the loan. If the company financed the project entirely with equity, the firm’s cost of capital would be 12 percent. The corporate tax rate is 30 percent. Using the adjusted present value method, calculate the APV of the project.

(Enter your answer in dollars, not millions of dollars, e.g., 1,234,567. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Calculation of net income for 4 years:

Earning before tax and depreciation = EBTD =  $608,000

As depreciation will be following straight line method, per year depreciation would be = $1,740,000 / 4 = $435000

Hence, Earning before tax and others = 608000 - 435000 = 173000

floating fee charged by the bank = 58000/ 4 = 14500

Earning before tax = 28000 (173000 - 145000)

As corporate tax is 30%, net income = 28000 * 0.7 = $19600

Operating Cash flow for each year = Net Income + Non cash expenses = 19600 + 435000 = 454600

Adjusted Present Value (APV) = CashFlow/(1 r) - Initial investment (where n = 4, r = cost of capital/ equity = 12%)

= 454600 (1/1.12 + 1/1.12^2 + 1/1.12^3 + 1/1.12^4) - 1,740,000 = 405892 + 362404 + 323575 + 288906 = 1,380,777 - 1740000 =  $-359222 (Answer)